Can you please help me solve: "A collection of dimes and nickels is worth $6.70. There are 92 coins in all. Find how many of each there are."
Answer by stanbon(75887)

Can you please help me solve: "A collection of dimes and nickels is worth $6.70. There are 92 coins in all. Find how many of each there are."
---------
Quantity Equation::: n + d = 92 coins
Value Equation:::: 5n +10d = 670 cents
------------------
Multiply thru the top equation by "t"
5n + 5d = 5*92
5n +10d = 670
------
Subtract the top from the bottom and solve for "d":
5d = 230
---
d = 46 (# of dimes)
----
Solve for "n":
n + d = 92
n + 46 = 92
n = 46 (# of nickels)
----------------------------
Cheers,
Stan H.